Hi, Im having problems trying to solve these rational equations. its a extraneous solution. Can you help please.
5/y-3 = y+7/2y-6 + 1

Answer by edjones(8007)

5/y-3 = y+7/2y-6 + 1
5/y-3 = y+7/2(y-3) + 1
10=y+7+2(y-3) Multiply each side of the equation by 2(y-3) to eliminate fractions.
y+7+2y-6=10
3y+1=10
3y=9
y=3
.
Now we replace y with 3 in the original equation and find that 5/y-3 becomes 5/3-3=5/0.
Since division by zero is not allowed then y=3 is an extraneous answer.
There is no answer.
.
Ed